OpenStudy (anonymous):

Solve. x2 + 10x = –2

OpenStudy (anonymous):

is that x^2?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

put the equation in the form:\[ax ^{2} + bx + c = 0\]and then solve using the quadratic formula:\[x = \frac{ -b \pm \sqrt{b ^{2} - 4ac} }{ 2a }\]Your equation re-written looks like: x^2 + 10x + 2 = 0

OpenStudy (anonymous):

My choices are: A. -5 +/- 2√2 B. -5 +/- 3√3 C. -5 +/- √7 D. -5 +/- √23

OpenStudy (anonymous):

So, a=1, b=10, and c=2. Substitute and simplify.

OpenStudy (anonymous):

i will try, but im doing it on paper because its easier to work out on there

OpenStudy (anonymous):

i got -10 +/- √92 all over 2

OpenStudy (anonymous):

You are on the right track. That accounts for substitution. Now just simplify.

OpenStudy (anonymous):

It's D

OpenStudy (anonymous):

because 4 is the greatest perfect square in 92

OpenStudy (anonymous):

If you don't show your work, then I can't tell you whether your answer is correct or not.

OpenStudy (anonymous):

92 divided by 4 is 23

OpenStudy (anonymous):

i did but it's on paper and will take a while to type out on here

OpenStudy (anonymous):

ok, yes, it's D.

OpenStudy (anonymous):

thanks for the help, i actually learned how to do it instead of just getting the answer

OpenStudy (anonymous):

That's the important thing. Because then, you'll be able to do these on your own.

OpenStudy (anonymous):

and you're very welcome. Thanks for the recognition and good luck to you @alassen